# Floquet-like theory and gauge transformations for general smooth dynamical systems

### v1

@@ Abstract
The classical Floquet theory allows to map a time-periodic system of linear differential equations into an autonomous one. By looking at it in a geometrical way, we extend the theory to a class of non-autonomous non-periodic equations. This is obtained by considering a change of variables which depends on time in a non-trivial way, i.e. introducing gauge transformations, well known in fundamental Physics and Field Theory – but which seems to have received little attention in this context.
@@ Acknowledgements
The work of GG is partially supported by the project “Mathematical Methods in Non-Linear Physics” (MMNLP) of INFN, and by GNFM-INdAM. In the course of this work GG enjoyed the warm hospitality and the relaxed work atmosphere of SMRI.
@@ 2 Floquet theory
This is completely general, i.e. it holds for whatever choice of the matrix $P(t)$ occurring in the change of variables.
@@
$F(y,t)\ =\ P^{-1}\,N(Ay,t)\ .$ (12)
Note that even in the case $\partial_{t}N(y,t)=0$, an explicit time-dependence will be introduced in $F$ both by $P^{-1}=P^{-1}(t)$ and by $A=A(t)$.
@@
Finally, note that examples can be built in which $F$ is actually independent of $t$: this is obtained e.g. by choosing $N(Ay,t)=P(t)N_{0}(Ay)$, see eq.(12). $\odot$
@@ 4.1 Analytical considerations
and we operate a gauge transformation (5) with a periodic matrix function $P(t)$. This will produce a non-autonomous linear system
@@
with $A(t)$ the periodic matrix function given precisely by (15). Actually, our previous discussion (that is, Floquet theorem) shows that any time-periodic linear system can be generated in this way from a constant coefficients system.
@@
Remark 3. We note that for any given $B$ and any initial condition for $P$, this system has a unique solution, as follows from the theory of linear differential equations; see [1] for details.
@@
the time dependence of ${\bf{\mathcal{F}}}$ originates in the time dependence of $P(t)$.
@@
In this case, in particular, if ${\bf F}$ is autonomous then also the new nonlinear term ${\bf\mathcal{F}}$ will be autonomous, and our autonomization of the time-dependent linear part will produce an overall autonomous system. $\odot$
@@ 5.1 Example 1
${\bf w}\ =\ R[\vartheta(t)]\ {\bf x}\ ,\ \ \ R[\vartheta]\ =\ \begin{pmatrix}\cos\vartheta&-\sin\vartheta\\ \sin\vartheta&\cos\vartheta\end{pmatrix}\ .$ (29)
@@ 6.1 Example 2
$\begin{cases}\xi^{\prime}\ =\ \omega(t)\,\eta\ +\ (1\,-\,\xi^{2}\,-\,\eta^{2})\,\xi&,\\ \eta^{\prime}\ =\ -\omega(t)\,\xi\ +\ (1\,-\,\xi^{2}\,-\,\eta^{2})\,\eta&.\end{cases}$ (34)
@@
$\begin{cases}\xi^{\prime}\ =\ \xi\ +\ \omega(t)\,\eta&,\\ \eta^{\prime}\ =\ -\omega(t)\,\xi\ +\ \eta&;\end{cases}$ (35)
@@
$A\ =\ \begin{pmatrix}1&\omega(t)\\ -\omega(t)&1\end{pmatrix}\ .$ (37)
@@
as follows from our general discussion. We should look for a $P=P(t)$ such that $B$ is a constant matrix. By definition the matrix $P$ satisfies
$P^{\prime}\ =\ A\ P\ -\ P\ B\ .$ (40)
@@
By (40), with $P$ of the form (42), the equation for $P$ reads as the system [^4] The matrix $P$ has four elements and hence the matrix equation for $P$ corresponds to a four dimensional system; however the special form of our initial equation (34) – note that $A(t)$ and $B$ commute – and our ansatz (42) for $P$ make that we have two pairs of identical equations.
$\begin{cases}\sin(\beta(t))\beta^{\prime}(t)\ =\ \omega(t)\sin(\beta(t))&,\\ \cos(\beta(t))\beta^{\prime}(t)\ =\ \cos(\beta(t))\omega(t)&.\end{cases}$ (43)
This is solved by $\beta^{\prime}(t)=\omega(t)$, i.e. by
@@
$\begin{cases}x^{\prime}\ =\ \left(1\ -\ (x^{2}+y^{2})\right)\ x&,\\ y^{\prime}\ =\ \left(1-(x^{2}+y^{2})\right)\ y&.\end{cases}$ (45)
@@ 6.2 Example 3
$\begin{cases}\xi^{\prime}\ =\ \omega(t)\,\eta\ +\ [1\,-\,R^{-1}(t)(\xi^{2}\,-\,\eta^{2})]\,\xi&,\\ \eta^{\prime}\ =\ -\omega(t)\,\xi\ +\ [1\,-\,R^{-1}(t)(\xi^{2}\,-\,\eta^{2})]\,\eta&.\end{cases}$ (46)
@@
$\begin{cases}x^{\prime}\ =\ x\ -\ R^{-1}(t)\ \left(x^{2}\,+\,y^{2}\right)\ x&,\\ y^{\prime}\ =\ y\ -\ R^{-1}(t)\ \left(x^{2}\,+\,y^{2}\right)\ y&.\end{cases}$ (47)
@@
[…]
@@ 6.3 Example 4
$\begin{cases}\xi^{\prime}\ =\ \omega(t)\,\eta\ +\ [1\,-\,\xi\,\eta]\,\xi&,\\ \eta^{\prime}\ =\ -\omega(t)\,\xi\ +\ [1\,-\,\xi\,\eta]\,\eta&.\end{cases}$ (48)
@@
$\begin{cases}x^{\prime}\ =\ x\ +\ \left[\frac{1}{2}\,\sin(2\beta(t))\,\left(y^{2}-x^{2}\right)\ -\ \cos(2\beta(t))\,x\,y\right]\ x&,\\ y^{\prime}\ =\ y\ +\ \left[\frac{1}{2}\,\sin(2\beta(t))\,\left(y^{2}-x^{2}\right)\ -\ \cos(2\beta(t))\,x\,y\right]\ y&.\end{cases}$ (49)
@@
here $I$ is the four-dimensional identity matrix, [.,.][.,.][ . , . ] is the commutator and $\epsilon_{ijk}$ is the completely alternating (Levi-Civita) symbol.
Given a generic matrix
$A\ =\ \sum_{i=1}^{3}\alpha_{i}\,Y_{i}\ ,\ \ \ \alpha_{1}^{2}+\alpha_{2}^{2}+\alpha_{3}^{2}=1\ ,$
@@ 7.1 Example 5
$K\ =\ \frac{\eta}{3-\cos^{2}(t)}\ \left(\begin{array}[]{llll}0&-1&\cos(\eta t)&\sin(\eta t)\\ 1&0&\sin(\eta t)&-\cos(\eta t)\\ -\cos(\eta t)&-\sin(\eta t)&0&-1\\ -\sin(\eta t)&\cos(\eta t)&1&0\end{array}\right)$ (53)
@@
$M\ =\ \frac{1}{\sqrt{2}}\ \left(\begin{array}[]{llll}\sin(t)&\cos(t)&\cos(\omega t)&\sin(\omega t)\\ -\cos(t)&\sin(t)&\sin(\omega t)&-\cos(\omega t)\\ -\cos(\omega t)&-\sin(\omega t)&\sin(t)&\cos(t)\\ -\sin(\omega t)&\cos(\omega t)&-\cos(t)&\sin(t)\end{array}\right)\ .$ (55)
@@ 7.2 Example 6
$K\ =\ \frac{1}{3-\cos^{2}t}\ \left(\begin{array}[]{llll}0&-a&(\eta+2)\cos(\eta t)&a\sin(\eta t)\\ a&0&\alpha\sin(\eta t)&-a\cos(\eta t)\\ -a\cos(\eta t)&-a\sin(\eta t)&0&-b\\ -a\sin(\eta t)&a\cos(\eta t)&b&0\end{array}\right)\ .$ (56)
@@ 8.1 Mapping to a linear system
$\frac{d}{dt}\begin{pmatrix}u\\ v\end{pmatrix}\ =\ \begin{pmatrix}g(t)+\alpha(t)&f(t)\\ -h(t)&\alpha(t)\end{pmatrix}\ \begin{pmatrix}u\\ v\end{pmatrix}\ ,$ (60)
@@
$A(t)\ =\ \begin{pmatrix}g(t)&f(t)\\ -h(t)&0\end{pmatrix}\ ,$ (61)
@@
Remark 8. It may be noted that a common time-dependent factor in the r.h.s. of (59) can be eliminated by a reparametrization of time; thus transformations multiplying the coefficients $f,g,h$ by a common factor should be seen as trivial, and one can restrict to consider $P:R\to SL(2,R)$. The connection between Riccati equations and gauge $SL(2,R)$ transformations has been widely studied, see e.g. [37, 38, 39, 40]. As far as we know, these studies did not put such use of gauge transformations into a general context for generic ODEs. $\odot$
@@ 8.2 Example 7
$\dot{{\bf w}}\ =\ A(t)\ {\bf w}\ ,\ \ A(t)\ =\ \begin{pmatrix}\kappa_{1}&\kappa_{0}\,e^{-\beta t}\\ -e^{-\beta t}&0\end{pmatrix}\ .$ (65)
@@
$B\ =\ \begin{pmatrix}\beta+\kappa_{1}&\kappa_{0}\\ 1&0\end{pmatrix}\ .$ (67)
@@ 8.3 Example 8
$P\ =\ \left(\begin{array}[]{ll}\frac{\tan(t)}{2}&-\sec(t)\\ \frac{1}{2}&0\end{array}\right)$ (72)
@@ 8.4 Example 9
This is singular in $t=0$ and in $t=-1$; we will hence think of this as being defined for $t>0$. (We could of course also consider this for $t<-1$, or for $-1<t<0$.)
@@
$A\ =\ \left(\begin{array}[]{ll}\frac{2-t^{2}}{t(t+1)}&\frac{2-t-t^{2}}{t^{2}(t+1)}\\ -(t+1)&0\end{array}\right)\ .$ (76)
@@
$P\ =\ \left(\begin{array}[]{ll}-\frac{t+1}{t}&-\frac{t+1}{t^{2}}\\ 1+\frac{1}{t}&\frac{1}{t^{2}}\end{array}\right)\ ,$ (77)
@@
$B\ =\ \begin{pmatrix}-1&1\\ 1&0\end{pmatrix}\ .$ (78)
@@ 9 Discussion and conclusions
Remark 10. Here we have been only concerned with dynamical systems; but it is clear that the method of considering gauge transformations to produce new systems amenable to some kind of exact treatment – and conversely, reduce new type of systems to those for which an exact treatment is available – is more general and not limited neither to dynamical systems neither to Floquet analysis; e.g. one might produce new integrable systems out of known ones.

### v3

# Open Communications in Nonlinear Mathematical Physics    ]ocnmp[   Vol.5 (2025) pp 1References   Article

@@ Abstract
The classical Floquet theory allows to map a time-periodic system of linear differential equations into an autonomous one. By looking at it in a geometrical way, we extend the theory to a class of non-autonomous non-periodic equations. This is obtained by considering a change of variables which depends on time in a non-trivial way, i.e. introducing gauge transformations, well known in fundamental Physics and Field Theory – but which seem to have received little attention in this context.
@@ Acknowledgements
We are grateful to an anonymous Referee for carefully checking our computations and for spotting several problemas with them. The work of GG is partially supported by the project “Mathematical Methods in Non-Linear Physics” (MMNLP) of INFN, and by GNFM-INdAM. In the course of this work GG enjoyed the warm hospitality and the relaxed work atmosphere of SMRI.
@@ 2 Floquet theory
This is completely general, i.e. it holds for whatever choice of the matrix $P(t)$ occurring in the change of variables. [^1] Hence the use of the notation $\widehat{A}$, leaving $B$ for the specific change of variables (4).
@@
$F(y,t)\ =\ P^{-1}\,N(Py,t)\ .$ (12)
Note that even in the case $\partial_{t}N(y,t)=0$, an explicit time-dependence will be introduced in $F$ by $P^{-1}=P^{-1}(t)$ and by $P=P(t)$.
@@
Finally, note that examples can be built in which $F$ is actually independent of $t$: this is obtained e.g. by choosing $N(x,t)=P(t)N_{0}(x)$, see eq.(12). $\odot$
@@ 4.1 Analytical considerations
and we operate a gauge transformation
x=Pyx\ =\ P\ y (18)
(note that here the role of $P$ is not the same as in (5), but is reversed) with a periodic matrix function $P(t)$. This will produce a non-autonomous linear system
@@
with $A(t)$ the periodic matrix function given by (15). Actually, our previous discussion (that is, Floquet theorem) shows that any time-periodic linear system can be generated in this way from a constant coefficients system.
@@
Remark 3. We note that for any given $B$ and any initial condition for $P$, this system has a unique solution, as follows from the theory of linear differential equations; see [1] for details. $\odot$
@@
[…]
@@
[…]
@@ 5.1 Example 1
${\bf w}\ =\ R[\vartheta(t)]\ {\bf x}\ ,\ \ \ R[\vartheta]\ =\ \begin{pmatrix}\cos\vartheta&-\sin\vartheta\\ \sin\vartheta&\cos\vartheta\end{pmatrix}\$. (30)
@@ 6.1 Example 2
[…]
@@
[…]
@@
$A\ =\ \begin{pmatrix}1&\omega(t)\\ -\omega(t)&1\end{pmatrix}\$. (38)
@@
as follows from our general discussion. We should look for a $P=P(t)$ such that $B$ is a constant matrix. By definition the matrix $P$ satisfies eq. (20), that is $P^{\prime}=AP-PB$.
@@
By (39), with $P$ of the form (42) and $B$ as prescribed by (41), the equation for $P$ reads as the system [^5] The matrix $P$ has four elements and hence the matrix equation for $P$ corresponds to a four dimensional system; however the special form of our initial equation (35) – note that $A(t)$ and $B$ commute – and our ansatz (42) for $P$ make that we have two pairs of identical equations.
@@
This reduces to $\beta^{\prime}(t)=-\,\omega(t)$, i.e. is solved by
$\beta(t)\ =\ \beta_{0}\ -\ \int_{0}^{t}\omega(\tau)\ d\tau\ .$ (44)
@@
[…]
@@ 6.2 Example 3
[…]
@@
[…]
@@
$\theta^{\prime}\ =\ 0\ ,\ \ \ \rho^{\prime}\ =\ \left(1\ -\ \frac{\rho^{2}}{R(t)}\right)\ \rho\ .$
@@ 6.3 Example 4
[…]
@@
[…]
@@
6.4 Example 5
The previous Examples are in a way abstract ones; we want now to discuss a specific one, i.e. an oscillator with a time dependent frequency; more specifically, we will consider the case where this frequency is quasi-periodic in time. The system we will consider can be integrated in an elementary way, so that the correctness of the result obtained by appying our approach can be checked directly.
Let $\Omega(t)$ denote the function
$\Omega(t)\ =\ \omega\ +\ a\,\sin(\alpha t)\ +\ b\,\sin(\beta t)\ ,$ (50)
where $a,b$ as well as $\omega,\alpha,\beta$ are nonzero real constants, and we will focus on the case where $\alpha/\beta$ is irrational; thus $\Omega(t)$ is a quasi-periodic function.
Consider the dynamical system
[…]
If $a$ and $b$ are small this represents a (quasi-periodic) perturbation of a harmonic oscillator with frequency $\omega$.
The system is promptly integrated by passing to polar coordinates $(\rho,\theta)$; in these we have $\rho=const$ and
$\theta(t)\ =\ \theta(0)\ +\ \int_{0}^{t}\Omega(\tau)\,d\tau\ .$ (52)
We want however, as mentioned above, to use this simple example to illustrate the application of our method in a concrete and physically meaningful case.
Our system (51) is written in vector notation as
$\dot{{\bf x}}\ =\ A(t)\ {\bf x}\ ,$ (53)
where ${\bf x}=(x,y)$ and
$A(t)\ =\ \Omega(t)\ J\ ,\ \ \ J\ =\ \begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\$. (54)
In order to handle the system with our method, we have to consider eq.(20), where both $P(t)$ and the constant matrix $B$ are unknown. We will make an ansatz for $B$ (see also Remark 3 in this respect), i.e. that it corresponds to the unperturbed problem obtained for $a=b=0$,
$B\ =\ \omega\ J\ .$
With this choice, our gauge-determining equation reduces to
$\frac{dP}{dt}\ =\ \Omega(t)\ J\,P\ -\ \omega\ P\,J\ .$ (55)
We will thus look for a solution in the form
$P(t)\ =\ \begin{pmatrix}\cos[\phi(t)]&-\sin[\phi(t)]\\ \sin[\phi(t)]&\cos[\phi(t)]\end{pmatrix}\$. (56)
With $P(t)$ of this form, the equation (55) reduces to
[…]
that is, $\phi(t)$ is determined by solving
$\phi^{\prime}(t)\ =\ \Omega(t)\ -\ \omega\ =\ A\,\sin(\alpha t)\ +\ B\,\sin(\beta t)\ .$ (58)
In conclusion, setting to zero the inessential integration constant, we get
$\phi(t)\ =\ -\,\frac{A}{\alpha}\,\cos(\alpha t)\ -\ \frac{B}{\beta}\,\cos(\beta t)\ ;$ (59)
plugging this into (56) we obtain the required gauge transformation. In the new variables
$\begin{pmatrix}\xi\\ \eta\end{pmatrix}\ =\ P\ \begin{pmatrix}x\\ y\end{pmatrix}$ (60)
the system reads just
[…]
We can them write the solution for these variables, and go back to the original ones by inverting (60); we obtain of course (52).
The computation is easily generalized to the case where the nonautonomous term is given by the superposition of $n$ irrational frequencies.
@@ 6.4 Example 5
$Y_{i}^{2}\ =\ -I\ ,\ \ Y_{i}\,Y_{j}\ =\ \epsilon_{ijk}\,Y_{k}\ ,\ \ [Y_{i},Y_{j}]\ =\ 2\,\epsilon_{ijk}\,Y_{k}\ ;$
here $I$ is the four-dimensional identity matrix, $[.,.]$ is the commutator and $\epsilon_{ijk}$ is the completely alternating (Levi-Civita) symbol.
Given a generic matrix
@@ 7.1 Example 6
[…]
@@
[…]
@@ 7.2 Example 7
[…]
@@ 8.1 Mapping to a linear system
$\frac{d}{dt}\begin{pmatrix}u\\ v\end{pmatrix}\ =\ \begin{pmatrix}g(t)+\alpha(t)&f(t)\\ -h(t)&\alpha(t)\end{pmatrix}\ \begin{pmatrix}u\\ v\end{pmatrix}\$, (72)
@@
$A(t)\ =\ \begin{pmatrix}g(t)&f(t)\\ -h(t)&0\end{pmatrix}\$, (73)
@@
Remark 8. It may be noted that a common time-dependent factor in the r.h.s. of (71) can be eliminated by a reparametrization of time; thus transformations multiplying the coefficients $f,g,h$ by a common factor should be seen as trivial, and one can restrict to consider $P:R\to SL(2,R)$. The connection between Riccati equations and gauge $SL(2,R)$ transformations has been widely studied, see e.g. [37, 38, 39, 40]. As far as we know, however, these studies did not put such a use of gauge transformations into a general context for generic ODEs. $\odot$
@@ 8.2 Example 8
$\dot{{\bf w}}\ =\ A(t)\ {\bf w}\ ,\ \ A(t)\ =\ \begin{pmatrix}\kappa_{1}&\kappa_{0}\,e^{-\beta t}\\ -e^{-\beta t}&0\end{pmatrix}\$. (77)
@@
$B\ =\ \begin{pmatrix}\beta+\kappa_{1}&\kappa_{0}\\ 1&0\end{pmatrix}\$. (79)
@@ 8.3 Example 9
[…]
@@ 8.4 Example 10
This is singular in $t=0$ and in t=1t=-1; we will hence think of this as being defined for $t>0$. (We could of course also consider this for t<1t<-1, or for 1<t<0-1<t<0.)
@@
[…]
@@
[…]
@@
$B\ =\ \begin{pmatrix}-1&1\\ 1&0\end{pmatrix}\$. (90)
@@ 9 Discussion and conclusions
Remark 10. Here we have been only concerned with dynamical systems; but it is clear that the method of considering gauge transformations to produce new systems amenable to some kind of exact treatment – and conversely, reduce new type of systems to those for which an exact treatment is available – is more general and not limited neither to dynamical systems nor to Floquet analysis; e.g. one might produce new integrable systems out of known ones.